OpenStudy (anonymous):

Use the sum or different to find he exact value of each trigonometric function a) tan(π)/(12)

OpenStudy (anonymous):

tan(pi/12)=tan(pi/6 - pi/12)

OpenStudy (anonymous):

a certain sense of deja vu

OpenStudy (anonymous):

what answer?

OpenStudy (anonymous):

tan pi/2=15 dgree

OpenStudy (anonymous):

tan15=tan(45-30)

OpenStudy (anonymous):

forget degrees, this has nothing to do with degrees

OpenStudy (anonymous):

i started to answer in last post, but only got half way, when mandolino linked to previous post here http://openstudy.com/users/mandolino#/users/mandolino/updates/4e86d5720b8b7ce881cf43a3 where it is worked out in all details

OpenStudy (anonymous):

\[\tan 15=\frac{\tan45-\tan30}{1-\tan45\tan30}\]

OpenStudy (anonymous):

"final answer" is \[2-\sqrt{3}\] but he or she wrote all the steps

OpenStudy (anonymous):

from here I have a trouble to solve

OpenStudy (anonymous):

\[2-\sqrt{3}\]is correct I don't know how get that answer

OpenStudy (anonymous):

you are confusing yourself by thinking you have to switch to degrees. but you have the formula backwards

OpenStudy (anonymous):

hold on let me find the right post

OpenStudy (anonymous):

thank you for the link I can find the link I post early

OpenStudy (anonymous):

\[\tan(\frac{\pi}{3}- \frac{\pi}{4})=\frac{\tan(\frac{\pi}{3})-\tan(\frac{\pi}{4})}{1-\tan(\frac{\pi}{3})\tan(\frac{\pi}{4})}\]

OpenStudy (anonymous):

so you need the following numbers to put in the formula \[\tan(\frac{\pi}{3})=\sqrt{3}\] and \[\tan(\frac{\pi}{4})=1\]

OpenStudy (anonymous):

put those numbers directly into the "subtraction angle" formula to get \[\tan(\frac{\pi}{12})=\frac{\sqrt{3}-1}{1-\sqrt{3}}\] and then rationalize the denominator to get your answer

OpenStudy (anonymous):

ty

OpenStudy (anonymous):

yw, hope it is clear